# Blackbox Dataset Inference for LLM

### v2

@@ 7. Related Works
Score-Based Membership Inference Attacks (MIAs) assume that the model behaves differently for training versus non-training data, which can be captured by some statistical metrics such as the perplexity (carlini2021extracting, ) or the loss values (yeom2018privacy, ). After gathering those metrics, they compare the model’s output score to the predefined threshold or use statistical methods to infer membership. Yeom et al. (yeom2018privacy, ) and Carlini et al. (carlini2021extracting, ) simply calculate the model’s loss or perplexity of given samples and subsequently apply a threshold to classify the samples as members or non-members. Mattern et al.  (mattern2023membership, ) propose the neighborhood attack, which generates highly similar ”neighbor” sentences using a pretrained masked language model and compares their losses to the original sample under the target model. If the original sample was part of the training data, the loss difference between it and its neighbors will be smaller than a predefined threshold, indicating membership. Min-kProb (shi2023detecting, ) calculate the average log likelihood of the ktokens with the lowest probabilities. A high average log-likelihood suggests the text is likely part of the pretraining data, as seen examples tend to lack outlier words with very low token probabilities. However, their effectiveness depends on the assumption that the model behaves differently for training versus non-training data, which may not hold if the model generalizes well.
@@
Reference Model-Based MIAs, on the other hand, involve training a separate reference model on a dataset similar to the target model’s training data. By comparing the behavior (e.g., perplexity or loss values) of the target model and the reference model, attackers can infer membership more robustly. Carlini et al. (carlini2021extracting, ) compares the perplexity ratio between a suspect model (e.g., a large GPT-2 model) and a smaller reference model (e.g., a smaller model) for a given text. If the suspect model has memorized the text during training, its perplexity will be significantly lower than the reference model’s, as the smaller reference model is not able to memorize as much training data as the suspect model.
Dataset Inference on LLM. Dataset Inference refers to the process of determining whether a specific dataset (or parts of it) was used to train a machine learning model. As far as we know, the research of Main et al.  (maini2024llm, ) is the only paper solving this problem in the LLM even in the natural language processing field. It points out that different MIAs perform well on specific datasets but poorly on others, and no single MIA consistently achieves high accuracy across all datasets. Therefore, they extract features from suspect and validation datasets using various MIAs, train a linear model to learn correlations between these features and dataset membership and perform a statistical T-Test to determine if the suspect dataset was used in training.
8. Conclusion

### v3

@@ VII Related Works
Score-Based Membership Inference Attacks (MIAs) assume models behave differently on training versus unseen data, which can be captured by some statistical metrics such as the perplexity [14] or the loss values [22]. After gathering those metrics, they compare the model’s output score to the predefined threshold or use statistical methods to infer membership. Mattern et al.  [7] propose the neighborhood attack, which generates highly similar ”neighbor” sentences using a pretrained masked language model and compares their losses to the original sample under the target model. If the original sample was part of the training data, the loss difference between it and its neighbors will be smaller than a predefined threshold, indicating membership. Min-k% Prob [6] calculate the average log likelihood of the k% tokens with the lowest probabilities. A high average log-likelihood suggests the text is likely part of the pretraining data, as seen examples tend to lack outlier words with very low token probabilities. However, their effectiveness depends on the assumption that the model behaves differently for training versus non-training data, which may not hold if the model generalizes well.
@@
Reference Model-Based MIAs, on the other hand, involve training a separate reference model on a dataset similar to the target model’s training data. By comparing the behavior (e.g., perplexity or loss values) of the target model and the reference model, attackers can infer membership more robustly. Carlini et al. [14] compares the perplexity ratio between a suspect model (e.g., a large GPT-2 model) and a smaller reference model (e.g., a smaller model) for a given text. If the suspect model has memorized the text during training, its perplexity will be significantly lower than the reference model’s, as the smaller reference model is not able to memorize as much training data as the suspect model.
Dataset Inference on LLM. Dataset inference determines whether a specific dataset was used in training. As far as we know, Maini et al. [9] present the only known work tackling this problem for LLMs. They show that different MIAs vary in performance across datasets, and no single attack performs consistently. Their approach extracts MIA-derived features from both suspect and validation datasets, trains a linear model to learn correlations, and applies a statistical T-Test to infer dataset inclusion.
VIII Conclusion